HUBBLE SPACE TELESCOPE DAILY REPORT #959 PERIOD COVERED: 0800 EST Daily Status Report as of 238/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make four standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. Two images were reported to be saturated, one was underexposed, and one was normal. 1.2 Completed HSP/POL 4887 (X-Ray Binaries -Aug.-Cycle 3 High) The High Speed Photometer was used to make at total of 36 photometric and polarimetric observations of X-Ray binary star LMCX-3 to obtain more information about the structure and condition in the gas streams connecting the primary star to its neutron companion. The observations were executed as scheduled, and all 36 observations were received and displayed. They all were reported to look okay. 1.3 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.4 Completed HSP/POL 5082 (X-Ray Binaries Part 2 - Cycle 3 High) The High Speed Photometer was used to make two photometric and polarimetric observations of X-Ray binary star HD153919 (aka, 3U1700-37) to obtain more information about the structure and condition in the gas streams connecting the primary star to its neutron companion. Both of the observations were received and displayed in real-time. The target was reported to be well centered. 1.5 Completed HSP/POL/UV2 5070 (Optical and Ultraviolet Observations of Radio Pulsars - cycle 2 High - Cont.) The High Speed Photometer was used to make observations of a pulsar which is located in the Large Magellanic Cloud. The UV2 detectors were used, and onboard target acquisition were used on a offset star. The Sample time was 0.0003 second, and the total observation time was about one hour. The OSS monitored the observations, and assessed the data quality. No problems were reported. 1.6 Completed WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make high latitude medium-deep sky survey observations taken in parallel with the prime science observations. The observations were received and displayed. They were reported to look okay. 1.7 Completed FGS 5060 (Linking HIPPARCOS to the Extragalactic Reference Frame - Part 2 of 6) The Fine Guidance Sensor were used to make astrometric observations of the object 2005-489IN. The observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. No problems were seen with the astrometric observations. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 12 Successful : 12 Losses of Lock : 1 Loss of Lock,from 237/2338Z to 2340Z The cause is still under investigation, subsequent reacquisition was successful. 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Significant HSTARs: None 2.4 Operations Note: The Joint Integration Simulation #1 (JIS #1, Rendezvous to Solar Array Stow and EVA #1) was successfully completed yesterday (Wednesday Aug. 25). The test lasted approximately 37 hours, and all major test objectives were met. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) 3.2 Complete Two Sets of HSP/POL/UV2 5082 (X-Ray binaries Part 2 - Cycle 3 High) 3.3 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.4 Complete FGS 5062 (Scale Cycle 2 Calibration: Fixed) 3.5 Complete FGS 5118 (Monitor the Long Term Stability of the FGSs Revised) 3.6 Complete Two Set of FGS 4892 (Duplicity Among Hyades Stars - Cycle 3 High) 3.7 Complete Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.8 Complete GHRS/2 4835 (Composition of Gas in Individual Interstellar Clouds: Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-